OpenStudy (anonymous):

y = ^3 square root 7x -9 help..i'm stumped :(

OpenStudy (anonymous):

What is it exactly that you want?

OpenStudy (anonymous):

do you mean the cube of square root of (7x-9)? y = (7x-8)^(3/2)

OpenStudy (anonymous):

I think he means \(y=\sqrt[3]{7x-9}.\)

OpenStudy (anonymous):

and as anwar said what do u want to do qith it?

OpenStudy (anonymous):

AnwarA has the correct equation. The question is what is the inverse of this equation. f^-1 = ??

OpenStudy (anonymous):

cube both sides: y^3 = 7x - 9 7x = y^3 + 9 x = (y^3 -9) / 7 replace y by x: f^-1 = (x^3 - 9) / 7